# Anomalous phonon Grüneisen parameters in semiconductor Ta2NiS5

### v1

@@ II Method
Strain-dependent Raman-scattering experiments were performed with a Horiba Jobin-Yvon LabRAM HR evolution spectrometer. We used a He-Ne laser (632.8 nm) with less than 1 mW power that was focused on the sample with a x50 magnification objective. The laser spot size was around 5 m in diameter. Spectra were recorded with 1800 mm-1 (0.6 cm-1 spectral resolution) and 600 mm-1 (1.6 cm-1 spectral resolution) gratings and a liquid-nitrogen-cooled CCD detector. All spectra of Raman response were corrected for the instrumental spectral response and Bose factor. Two polarization configurations were employed to probe excitations in different symmetry channels, namely the and configurations in which the incident light polarization lies along the crystallographic direction, and scattered one along or direction, respectively. According to Raman selection rules [33], in the orthorhombic phase (point group D2h), the and geometries probe the Ag and B2g Raman signal, respectively. Samples with a cleaved crystallographic plane were glued to Razorbill CS220T strain cell by Loctite Stycast 2850FT epoxy with CAT 24LV as the epoxy catalyst. The sample surfaces were examined under polarized light to find an area without intrinsic strain. We verified that the process of mounting the sample onto the strain cell does not introduce additional stress by comparing the Raman spectra of unmounted and mounted samples. The strain transmission was estimated from XRD measurement as detailed in the supplemental material [32]; the corresponding data measured at European Synchrotron Radiation Facility (ESRF) ID15B beamline are available from the ESRF Data Portal [34].
@@ III.1 Strain-free phonon spectra
The phonon leakage is an indication of symmetry breaking, and the phonon softening suggests a tendency towards an orthorhombic-to-monoclinic structural distortion. A recent XRD experiment finds no clear evidence for a structural transition [22]. Our in-house XRD investigation could not reveal any signature of such transition down to 80 K [32]. We note that a refinement of the crystal structure of Ta2NiS5 at 15 K with data taken at the synchrotron is not yet fully conclusive and calls for additional studies to clarify this issue [32].
Nonetheless, another noteworthy feature of the Raman spectra of Ta2NiS5 is the anomalous softening of the two lowest-energy B2g modes (hereafter labelled B and B) on cooling. As previously noted [20], this behavior is inconsistent with the conventional anharmonic decay of phonons in solids and might indicate the proximity of a phase transition. Because of the lattice contraction upon cooling, the corresponding Grüneisen parameter of these phonons is negative, which indicates that the orthorhombic structure is generally unstable against a reduction of the unit cell volume. Interestingly, the order parameter of the orthorhombic-to-monoclinic transition in Ta2NiSe5 has the B2g symmetry and the vibration pattern of the two anomalous modes consists in a shear displacement of the Ta along the a-direction, reminiscent of the structural distortion (in contrast with the B mode which involves mostly S vibrations and behaves normally). Taken together, these observations indicate that lattice strain tuning might be a promising tuning knob of the electronic properties of this system, as discussed below.
@@ III.3 First-principles calculations and discussion
We conclude from this that the anomalous behavior of the B phonon, both as function of temperature and of strain, is rooted in an incipient instability of the orthorhombic lattice structure against a shear monoclinic distortion of the Ta chains, which is likely to occur irrespective of the electronic gap size. Additional work is required to demonstrate whether such this distortion is effectively realized in Ta2NiS5, as suggested by the Raman data, but our work demonstrates that the susceptibility towards it is very large, and would likely occur upon compression of the lattice. At the same time, such perturbation would allow a reduction of the semiconducting gap size, which would offer an interesting way to approach the regime in which an EI can occur (i.e. when the gap size becomes lower than the exciton binding energy).

### v2

@@ II Method
Strain-dependent Raman-scattering experiments were performed with a Horiba Jobin-Yvon LabRAM HR evolution spectrometer. We used a He-Ne laser (632.8 nm) with less than 1 mW power that was focused on the sample with a x50 magnification objective. The laser spot size was around 5 m in diameter. One notch filter and two Bragg filters were used in the collection optical path to filter the elastically scattered light. Spectra were recorded with 1800 mm-1 (0.6 cm-1 spectral resolution) and 600 mm-1 (1.6 cm-1 spectral resolution) gratings and a liquid-nitrogen-cooled CCD detector. All spectra of Raman response were corrected for the instrumental spectral response and Bose factor. Two polarization configurations were employed to probe excitations in different symmetry channels, namely the and configurations in which the incident light polarization lies along the crystallographic direction, and scattered one along or direction, respectively. According to Raman selection rules [33], in the orthorhombic phase (point group D2h), the and geometries probe the Ag and B2g Raman signal, respectively. Samples with a cleaved crystallographic plane were glued to Razorbill CS220T strain cell by Loctite Stycast 2850FT epoxy with CAT 24LV as the epoxy catalyst. The sample surfaces were examined under polarized light to find an area without intrinsic strain. We verified that the process of mounting the sample onto the strain cell does not introduce additional stress by comparing the Raman spectra of unmounted and mounted samples. The strain transmission was estimated from XRD measurement as detailed in the supplemental material [32]; the corresponding data measured at European Synchrotron Radiation Facility (ESRF) ID15B beamline are available from the ESRF Data Portal [34].
@@ III.1 Strain-free phonon spectra
The phonon leakage is an indication of symmetry breaking, and the phonon softening suggests a tendency towards an orthorhombic-to-monoclinic structural distortion. A recent XRD experiment finds no clear evidence for a structural transition [22]. Our in-house XRD investigation could not reveal any signature of such transition down to 80 K [32] and our refinement of the crystal structure of Ta2NiS5 at 15 K with data taken at the synchrotron is not yet fully conclusive [32]. As a result additional studies are called for to clarify this issue.
Nonetheless, another noteworthy feature of the Raman spectra of Ta2NiS5 is the anomalous softening of the two lowest-energy B2g modes (hereafter labelled B and B) on cooling. As previously noted [20], this behavior is inconsistent with the conventional anharmonic decay of phonons in solids and might indicate the proximity of a phase transition. Because of the lattice contraction upon cooling, the corresponding Grüneisen parameter of these phonons is negative, which indicates that the orthorhombic structure is generally unstable against a reduction of the unit cell volume. Interestingly, the order parameter of the orthorhombic-to-monoclinic transition in Ta2NiSe5 has the B2g symmetry and the vibration pattern of the two lowest-energy B2g modes, which couple to the overdamped excitonic mode [19], consists of a shear displacement of the Ta along the a-direction, reminiscent of the structural distortion (in contrast with the B mode which involves mostly S vibrations and behaves normally). Nevertheless, the relevant B2g modes in Ta2NiSe5 exhibit no softening behavior [19, 20], consistent with the view that the transition is driven by an excitonic instability. Taken together, these observations indicate that lattice strain tuning might be a promising tuning knob of the electronic properties of this system, as discussed below.
@@ III.3 First-principles calculations and discussion
From these discussions we see that the anomalous behavior of the B phonon, both as function of temperature and of strain, is rooted in an incipient instability of the orthorhombic lattice structure against a shear monoclinic distortion of the Ta chains, which is likely to occur irrespective of the electronic gap size. Time-resolved spectroscopy, which has been successfully used in the study of Ta2NiSe5 to disentangle the charge and lattice dynamics [46, 47, 48, 16, 17], would provide further support to this view.
Although additional work is required to examine whether this distortion is effectively realized in Ta2NiS5, as suggested by the Raman data, our work demonstrates that the susceptibility towards it is very large, and would likely occur upon compression of the lattice. At the same time, such a perturbation would allow a reduction of the semiconducting gap size, which would offer an interesting way to approach the regime in which an EI can occur (i.e. when the gap size becomes lower than the exciton binding energy).
@@ Acknowledgements.
- Mor et al. [2017] S. Mor, M. Herzog, D. Golež, P. Werner, M. Eckstein, N. Katayama, M. Nohara, H. Takagi, T. Mizokawa, C. Monney, and J. Stähler, Ultrafast electronic band gap control in an excitonic insulator, Phys. Rev. Lett. 119, 086401 (2017).
- Mor et al. [2018] S. Mor, M. Herzog, J. Noack, N. Katayama, M. Nohara, H. Takagi, A. Trunschke, T. Mizokawa, C. Monney, and J. Stähler, Inhibition of the photoinduced structural phase transition in the excitonic insulator , Phys. Rev. B 97, 115154 (2018).
- Bretscher et al. [2021] H. M. Bretscher, P. Andrich, P. Telang, A. Singh, L. Harnagea, A. K. Sood, and A. Rao, Ultrafast melting and recovery of collective order in the excitonic insulator ta2nise5, Nature Communications 12, 1699 (2021).